OpenStudy (anonymous):

Express as a single log and simplify, if possible: logb (b is the subscript) 1. logb (x ^3 + y^3) - logb (x+y) = Expand: 1. logb (x+1)^3 / y^3 = PLEASE HELP!

OpenStudy (anonymous):

\[\log _{b} (x ^3 + y^3) - \log _{b} (x+y) \]

hartnn (hartnn):

log A + log B = log AB

hartnn (hartnn):

log A - log B = log A/B

OpenStudy (anonymous):

\[\frac{ \log_{a} (x + 1) ^{3} + x }{ y^{3} }\]

hartnn (hartnn):

log A - log B = log A/B so whats logb (x ^3 + y^3) - logb (x+y) =.. ?

OpenStudy (anonymous):

@hartnn yes, we just had it in class but i don't know what to do with this one. :(

OpenStudy (anonymous):

\[\frac{ logb(x+y)(x^2-xy+y^2) }{logb(x+y) }\]

OpenStudy (anonymous):

can you do from there?

OpenStudy (anonymous):

@hartnn if i have to factor it out, i'd get \[ \log _{b} x^{2} - xy + y^{2} \]

OpenStudy (anonymous):

@winterfez yes. and then what? i'm stuck right there. :)

hartnn (hartnn):

 did u get this ?